A 50 Hz sinusoidal voltage has rms value of 200 V. At t=0, the instantaneous value is positive and half of its maximum value. Write down expression for voltage and sketch the waveform


1
Expert's answer
2022-01-30T13:51:29-0500

Using the provided data, we proceed to find the elements for "V_{(t)}=V_{max}\sin{(\omega t+\phi)}":


"\\ f = 50\, Hz 
\\ \therefore \omega=2\pi f= 314.16\,{ rad/s}"


We find the maximum value with Vrms and the phase angle as:


"V_{rms}=200\,V
\\V_{max}=\sqrt{2}V_{rms}= (1.4142)(200\,V)=282.84\,V
\\ \phi= \arcsin{\Big( \dfrac {V_{(0)}}{V_{max}}\Big)} 
\\ \text{}
\\ \phi= \arcsin{\Big( \dfrac {½V_{max}}{V_{max}}\Big)} =\arcsin{(½)}= 30°=\frac{\pi}{6}"


In conclusion, the expression for the voltage will be "V_{(t)}=(282.84\,V)\sin{(100\pi t+\frac{\pi}{6})}".


The graphic for the voltage will be: